OpenStudy (anonymous):

for the given equation list the intercepts and test for symmetry y=-x^7/x^6-8

OpenStudy (anonymous):

\[y =\frac{ -x^7 }{ x^6-8 },\frac{ -x^7 }{ x^6 }-8\]which one

OpenStudy (anonymous):

\[0=\frac{ -x^7 }{ x^6-8 }\] \[x^7=0\] \[x=0\] \[y=0\] intercetpts\[(0,0)(0,0)\]

OpenStudy (anonymous):

So, the intercepts are (0,0)(0,0)...how do I test for symmetry?

OpenStudy (anonymous):

see is this is true f(x)=f(-x)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Actually, since f(x)=f(-x) it is not true, right?

OpenStudy (anonymous):

yes its not true hence the function is odd ie not symmetrical

OpenStudy (anonymous):

Thank you so much for your help!